What is the ratio of escape speed of the earth and escape speed of the moon?

Grade:11

1 Answers

Arun
6 years ago
Dear student

Escape velocity is given by,
v = \sqrt(2G*m/r) = \sqrt(2*G*r)
V(earth)/V(moon)  = \sqrt[g(earth)*r(earth)] /\sqrt[g(moon)*r(moon)]
Since g(earth) = 6* g(moon)
r(earth) = 4* r(moon)
When put these values
V(earth)/V(moon) = \sqrt24 /1
Regards
Arun (askIITians forum expert)